OpenStudy (anonymous):

There are 135 girls in a club of 415 students. Find the ratio of girls to boys in the club.

OpenStudy (anonymous):

415-135 = no. of boys = 280 no of girls/ no of boys = 135/280

OpenStudy (anonymous):

g+b=415->b=415-135->g/b=135/415-135

OpenStudy (anonymous):

answet choices 13:41 27:56 3:17 1:3

OpenStudy (anonymous):

27:56

OpenStudy (anonymous):

as simplify 135/280 by diving top and bottom by 5